State the order of reaction with respect to a reactant, R, if:
a). the rate of reaction doubles when [R] doubles
b). the rate of reaction is unchanged when [R] triples
c). the rate of reaction reduces by a factor of four when [R] is halved.
1
Expert's answer
2020-11-02T08:47:37-0500

State the order of reaction with respect to a reactant, R, if:


a) the rate of reaction doubles when [R] doubles


b) the rate of reaction is unchanged when [R] triples


c) the rate of reaction reduces by a factor of four when [R] is halved

Solution

 "v_R=-d[R]\/dt=k[R]^n" - rate equation [1]

a) "v_1=v_2\/2" and "[R_1]=[R_2]\/2"

Hense 

"k[R_1]^n=k[R_2]^n\/2"

"([R_1]\/[R_2])^n=1\/2"

"n=log_{[R_1] \above{2pt} [R_2]}0,5= log_{1 \above{2pt} 2}0,5=1"

b) "v_1=v_2" and "[R_1]=[R_2]\/3"

Hence

"k[R_1]^n=k[R_2]^n"

"([R_1]\/[R_2])^n=1"

"n=log_{[R_1] \above{2pt} [R_2]}1=0"

c)"v_1=4v_2" and "[R_1]=2[R_2]"

Hence

"k[R_1]^n=4k[R_2]^n"

"([R_1]\/[R_2])^n=4"

"n=log_{[R_1] \above{2pt} [R_2]}4= log_{2}4=2"

Answer: a) "n=1", b) "n=0", c) "n=2"

Reference: https://en.wikipedia.org/wiki/Rate_equation